# On Modular Invariants of Truncated Polynomial Rings in low ranks

### v1

@@ 2.2. The $\Delta$ subspace of the Dickson algebra
Given a positive integer $s$ and a partition $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{s}$ with at most $s$ nonzero parts, we say that a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in ${\mathcal{D}}_{s}$ is of type $(\lambda_{1},\ldots,\lambda_{s})$ (or q-compatible with $(\lambda_{1},\ldots,\lambda_{s})$, in the language of [15, Definition 5.4]) if for each $1\leq i\leq s$,
@@
It is clear that each Dickson monomial is $q$-compatible with a unique partition $(\lambda_{1},\ldots,\lambda_{s})$ of length $\leq s$.
Now if $s\leq m$ then we restrict our attention to partitions which fit inside a rectangular $s\times(m-s)$, that is partitions with at most $s$ nonzero parts and $m-s\geq\lambda_{1}$. Define $\Delta_{(\lambda_{1},\ldots,\lambda_{s})}$ as the subset of the Dickson algebra ${\mathcal{D}}_{s}$ consisting of Dickson monomials of type $(\lambda_{1},\ldots,\lambda_{s})$, and let $\Delta^{m}_{s}$ denote the disjoint union of all subsets over all such partitions fitting inside $s\times(m-s)$. By convention, we put $\Delta^{m}_{s}=\emptyset$ if $s>m$, $\Delta^{m}_{0}=\{1\}$.
The Dickson monomials in $\Delta^{m}_{s}$ for $s\leq\min(m,n)$ together with the delta operators recalled below are the building blocks for our basis. For this reason, we will say that a Dickson monomial in $\Delta^{m}_{s}$ an essential monomial. We will also work with the edge monomials, which is a Dickson monomial not in $\Delta^{m}_{s}$ but has the form $Q_{s,i}f$ where $f\in\Delta^{m}_{s}$, for some $10\leq i\leq s-$ 1. Clearly, the edge monomials are those such that $e_{i}\in[\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1})$ for all $i\neq j$ and $e_{j}=\frac{q^{\lambda_{j}+1}-q^{\lambda_{j+1}}}{q-1}$ for some $1\leq j\leq s$.
It is easy to check that $\Phi\Delta^{m}_{s}\subset\Delta^{m}_{s+1}$. In fact, $\Phi$ sends a Dickson monomial in ${\mathcal{D}}_{s}$ of type $(\lambda_{1},\ldots,\lambda_{s})$ to a Dickson monomial in ${\mathcal{D}}_{s+1}$ of type $(\lambda_{1},\ldots,\lambda_{s},0)$.
@@ 2.3. The delta operator
In this paper, with the number $m$ fixed from the beginning, we will always write $\delta_{s}$ for $\delta_{s;m}$ and occasionally use the full notation $\delta_{a;b}$ when necessary.
It was shown in greater generality in [6] that $\delta_{s}(f)$ is generally not a polynomial, but it is very often a polynomial and is a $\operatorname{GL}_{s}$-invariant. We give a precise statement and provide an elementary proof in the case for $\delta_{3}$ to make the paper self-contained.
@@
If $f$ is a $\operatorname{GL}_{2}$-invariant polynomial, then $\delta_{3}(f)$ is a polynomial and is $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
@@
We first show that $\delta_{3}(f)$ is a polynomial. The denominator $L_{3}$ is a product of linear forms, so it suffices to prove that the numerator becomes zero whenever there exists a nontrivial linear relation $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$. This is reduced to the proof that
$a_{1}x_{1}^{q^{m}}f(x_{2},x_{3})+a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=0,$
provided $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$ and not all scalars $a_{i}$ are zero. Without loss of generality, assume that $a_{3}=-1$. If both $a_{1},a_{2}\neq 0$, then since $f$ is a $\operatorname{GL}_{2}$-invariant, we have
@@
and the above expression becomes
$[a_{1}x_{1}^{q^{m}}+a_{2}x_{2}^{q^{m}}+a_{3}x_{3}^{q^{m}}]f(x_{1},x_{2})=[a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}]^{q^{m}}f(x_{1},x_{2})=0.$
If one of the $a_{i}$ is zero, say $a_{1}=0$, then $f(x_{2},x_{3})=f(x_{2},a_{3}x_{2})=0$ and $f(x_{1},x_{3})=f(x_{1},x_{2})$, so that the sum is reduced to
$a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=(a_{2}x_{2}+a_{3}x_{3})^{q^{m}}f(x_{1},x_{2})=0.$
If $a_{1}=a_{2}=0$ or $x_{3}=0$, this is also clear. We have proved that $\delta_{3}(f)$ is a polynomial.
Next we show that $\delta_{3}(f)$ is $\operatorname{GL}_{3}$-invariant modulo $I_{m}$. Clearly, it is symmetric, so it suffices to prove that $\delta_{3}(f)$ does not change under the operation that sends $x_{1}\mapsto x_{1}+x_{2}$ and fixes $x_{2},x_{3}$. The denominator $L_{3}$ certainly satisfies this condition. By direct inspection, $\delta_{3}(f)(x_{1}+x_{2},x_{2},x_{3})-\delta_{3}(f)(x_{1},x_{2},x_{3})$ equals
@@
For this expression to be zero in ${\mathcal{Q}}_{m}(3)$, we must prove that the second factor is a polynomial. The rest of the argument is similar to the first part of the proof. ∎
@@
The $\operatorname{GL}_{1}$-polynomials are of the form $f(x)=x^{s(q-1)}=Q_{1,0}^{s}$, $s\geq 0$. Then $\delta_{2}(x^{s(q-1)})$ is a $\operatorname{GL}_{2}$-invariant polynomial in ${\mathcal{Q}}_{m}(2)$. For $s<[m]_{q}$, this is Goyal’s $y_{s}$ family in [5] where it was defined as
@@
Our construction is also valid for $s\geq[m]_{q}$. However, it is trivial in ${\mathcal{Q}}_{m}$ except when $s=[m]_{q}+1$, in which case
@@
It is easy to check that $y_{s}$ is a genuine $\operatorname{GL}_{2}$-invariant (that is, a Dickson polynomial) whenever $s=\frac{q^{i}-1}{q-1}$ for some non-negative integer $i$.
@@
For $s\leq[m]_{q}$, $\delta_{2}^{2}(Q_{1,0}^{s})$ is a $\operatorname{GL}_{3}$-invariant of degree $2(q^{m}-q)+s(q-1)$ in ${\mathcal{Q}}_{m}$. This family of elements coincides with Goyal’s polynomials $a_{m,3,s}$ in [5, Corollary 4.3] where it was written in a different form:
@@
where the sum runs over all triple $(i_{1},i_{2},i_{3})$ of non-negative integers such that $i_{s}<[m]_{q}$ and $i_{1}+i_{2}+i_{3}=2[m]_{q}-2+s$. Again, our definition works also for $s>[m]_{q}$:
$\delta_{2}^{2}(Q_{1,0}^{[m]_{q}+2})=\delta_{1}^{3}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$
@@ 2.4. The delta operator and the Dickson algebra
For each composition $\alpha$ of $n$, the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is a module over the Dickson algebra ${\mathcal{D}}_{n}$. The next result describes how the delta operator interacts with the Dickson algebra in low rank.
@@
We will prove only statements (3) for $i=1$ and the last two parts, as the proof of other statements use similar arguments and are even more straightforward.
@@
Proof of (3). Consider the difference $Q_{3,1}\delta_{3}(f)-\delta_{3}(Q_{2,0}^{q}f)$. Using the identity $Q_{3,1}=V_{3}^{q-1}Q_{2,1}+Q_{2,0}^{q}$, the (3,1)-entry in the last row of the determinant in the numerator can be simplified as
$x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3}).$
@@
Since $V_{3}(x_{2},x_{3},x_{1})L_{2}(x_{2},x_{3})=L_{3}(x_{2},x_{3},x_{1})=L_{3}(x_{1},x_{2},x_{3})$, the Laplace expansion along the last row of the determinant in the numerator shows that the resulting polynomial belongs to $I_{m}$, which is zero in ${\mathcal{Q}}_{m}$.
Proof of (4). Write $Q_{3,2}=V_{3}^{q-1}+Q_{2,1}^{q}$. From (1) and (2) we have in ${\mathcal{Q}}_{m}$
$Q_{3,2}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,1}^{q}\delta_{2}f).$
Using the formula $Q_{2,1}=V_{2}^{(q-1)}+Q_{1,0}^{q}$ and unravelling the definition, we see that the difference $\delta_{2}(Q_{2,1}^{q}\delta_{2}f)-\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ can be written as the sum
$x_{3}^{q^{m}}\frac{x_{1}^{q^{m}}x_{2}f(x_{2})V_{2}^{q^{2}-q-1}(x_{2},x_{3})-x_{2}^{q^{m}}x_{1}f(x_{1})V_{2}^{q^{2}-q-1}(x_{1},x_{3})}{L_{2}(x_{1},x_{2})}+\\ x_{1}^{q^{m}}x_{2}^{q^{m}}f(x_{3})\frac{x_{2}V_{2}^{q^{2}-q-1}(x_{3},x_{2})-x_{1}V_{2}^{q^{2}-q-1}(x_{3},x_{1})}{L_{2}(x_{1},x_{2})}.$
It is clear that the first summand is a polynomial, and is a multiple of $x_{3}^{q^{m}}$. The second summand belongs to the ideal $(x_{1}^{q^{m}},x_{2}^{q^{m}})$.
Proof of (5). We have in ${\mathcal{Q}}_{m}$
$Q_{3,1}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,0}^{q}\delta_{2}f)$
and
$Q_{2,0}^{q}\delta_{2}(f)=L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{1}f(x_{1})x_{2}^{q^{m}}-L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{2}f(x_{2})x_{1}^{q^{m}}.$
The rest of the argument is similar to (4) above. ∎

### v2

@@ 2.2. The $\Delta$ subspace of the Dickson algebra
Lewis, Reiner and Stanton [9, Subsection 7.4] speculated that a portion of the Dickson algebra studied in [12, Section 5] might form part of an $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}(n)^{G}$. Our work confirms this intuition. In fact, this portion, which we call the Delta subspace, plays a central role. We now review its construction and some elementary properties.
Given a positive integer $s$ and a partition $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{s}$ with at most $s$ nonzero parts, we say that a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in ${\mathcal{D}}_{s}$ is of type $(\lambda_{1},\ldots,\lambda_{s})$ (or q-compatible with $(\lambda_{1},\ldots,\lambda_{s})$, in the language of [12, Definition 5.4]) if for each $1\leq i\leq s$,
@@
By convention, $\lambda_{s+1}=0$. Every Dickson monomial in ${\mathcal{D}}_{s}$ is $q$-compatible with a unique partition $(\lambda_{1},\ldots,\lambda_{s})$ of length $\leq s$.
Now suppose $s\leq m$, and we will restrict our attention to partitions whose Ferrers diagrams fit inside a $s\times(m-s)$ rectangle, that is, partitions with at most $s$ nonzero parts and $m-s\geq\lambda_{1}$. Define $\Delta_{(\lambda_{1},\ldots,\lambda_{s})}$ as the subset of the Dickson algebra ${\mathcal{D}}_{s}$ consisting of Dickson monomials of type $(\lambda_{1},\ldots,\lambda_{s})$, and let $\Delta^{m}_{s}$ denote the disjoint union of all subsets over all such partitions fitting inside an $s\times(m-s)$ rectangle. By convention, we put $\Delta^{m}_{s}=\emptyset$ if $s>m$, and $\Delta^{m}_{0}=\{1\}$.
The Dickson monomials in $\Delta^{m}_{s}$ for $s\leq\min(m,n)$ together with the delta operators recalled below are the building blocks for our basis. For this reason, we will say that Dickson monomials in $\Delta^{m}_{s}$ essential monomials. We will also work with the edge monomials, which are the Dickson monomials not in $\Delta^{m}_{s}$ that have the form $Q_{s,i}f$ where $f\in\Delta^{m}_{s}$, for some $10\leq i\leq s-$ 1. Clearly, the edge monomials are those such that $e_{j}=\frac{q^{\lambda_{j}+1}-q^{\lambda_{j+1}}}{q-1}$ for some $1\leq j\leq s$ and $e_{i}\in[\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1})$ for all $i\neq j$.
@@ 2.3. The delta operator
It was shown in greater generality in [6] that $\delta_{s}(f)$ is generally not a polynomial. Still, in the cases that we are interested in, it is a genuine polynomial. Moreover, $\delta_{s}(f)$ is rarely a Dickson polynomial but it is a $G_{s}$-invariant modulo $I_{m}$. We give a precise statement and provide an elementary proof for $\delta_{3}$ to make the paper self-contained.
@@
If $f$ is a $G_{2}$-invariant polynomial, then $\delta_{3}(f)$ is also a polynomial and is $G_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
@@
We first show that $\delta_{3}(f)$ is a polynomial. Recall the denominator $L_{3}=V_{1}V_{2}V_{3}$ is a product of nonzero linear forms in the variables $x_{1},x_{2},x_{3}$, so it suffices to prove that the numerator of $\delta_{3}(f)$ becomes zero whenever there exists a nontrivial linear relation among the $x_{i}$, which by symmetry, can be assumed to have the following form:
(2.1) $x_{3}=a_{1}x_{1}+a_{2}x_{2},\quad(a_{1},a_{2})\in\operatorname{\mathbb{F}}_{q}\times\operatorname{\mathbb{F}}_{q}.$
We claim that this implies a similar linear relation among the columns of the numerator of $\delta_{3}(f)$, that is,
(2.2) $x_{3}^{q^{m}}f(x_{1},x_{2})=a_{1}x_{1}^{q^{m}}f(x_{2},x_{3})+a_{2}x_{2}^{q^{m}}f(x_{1},x_{3}).$
Indeed, this is clear if $a_{1}=a_{2}=0$. If the scalars $a_{1},a_{2}$ are both nonzero, then there is an equality
@@
since $f$ is a $G_{2}$-invariant. So the right-hand side of (2.2) becomes
$(a_{1}x_{1}^{q^{m}}+a_{2}x_{2}^{q^{m}})f(x_{1},x_{2})=(a_{1}x_{1}+a_{2}x_{2})^{q^{m}}f(x_{1},x_{2})=x_{3}^{q^{m}}f(x_{1},x_{2}).$
Finally, if one of the $a_{i}$ is zero, say $a_{1}=0$, Equation (2.2) is reduced to
$(a_{2}x_{2})^{q^{m}}f(x_{1},x_{2})=a_{2}x_{2}^{q^{m}}f(x_{1},a_{2}x_{2}),$
which is again true because $f$ is $G_{2}$-invariant. We have finished the proof that $\delta_{3}(f)$ is a polynomial.
The next step is to show that $\delta_{3}(f)$ is $G_{3}$-invariant modulo $I_{m}$. By symmetry, it suffices to prove that $\delta_{3}(f)$ does not change under the operation that sends $x_{1}$ to $x_{1}+x_{2}$ and fixes $x_{2},x_{3}$. The denominator $L_{3}$ certainly satisfies this condition. By direct inspection, the difference $\delta_{3}(f)(x_{1}+x_{2},x_{2},x_{3})-\delta_{3}(f)(x_{1},x_{2},x_{3})$ equals
@@
We already know that this is a polynomial. Moreover, $x_{2}$ occurs as a simple factor in the denominator $L_{3}$, and the determinant in the numerator vanishes when $x_{2}=0$. It follows that this polynomial is divisible by $x_{2}^{q^{m}}$, thus trivial in ${\mathcal{Q}}_{m}(3)$. ∎
@@
The $G_{1}$-polynomials are of the form $f(x)=x^{s(q-1)}=Q_{1,0}^{s}$, $s\geq 0$. Then $\delta_{2}(x^{s(q-1)})$ is a $G_{2}$-invariant polynomial in ${\mathcal{Q}}_{m}(2)$. For $s<[m]_{q}$, this is Goyal’s $y_{s}$ family in [5] where it was defined as
@@
Our construction is also valid for $s\geq[m]_{q}$. However, it is trivial in ${\mathcal{Q}}_{m}(2)$ except when $s=[m]_{q}+1$, in which case
@@
It is easy to check that $y_{s}$ is a genuine $G_{2}$-invariant (that is, a Dickson polynomial) whenever $s=\frac{q^{i}-1}{q-1}$ for some non-negative integer $i$.
@@
For $s\leq[m]_{q}$, $\delta_{2}^{2}(Q_{1,0}^{s})$ is a $G_{3}$-invariant of degree $2(q^{m}-q)+s(q-1)$ in ${\mathcal{Q}}_{m}(3)$. This family of elements coincides with Goyal’s polynomials $a_{m,3,s}$ in [5, Corollary 4.3] where it was written in a different form:
@@
where the sum is taken over all triples $(i_{1},i_{2},i_{3})$ of non-negative integers such that $i_{s}<[m]_{q}$ and $i_{1}+i_{2}+i_{3}=2[m]_{q}-2+s$. Again, our definition also works for $s>[m]_{q}$:
$\delta_{2}^{2}(Q_{1,0}^{[m]_{q}+2})=\delta_{1}^{3}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1},$
@@
Warning 2.6.
It is not true in general that if $f=g$ in ${\mathcal{Q}}_{m}(n)$ then $\delta_{s}(f)=\delta_{s}(g)$ in ${\mathcal{Q}}_{m}(n)$, even in the case where all expressions involved are polynomials. This is evident from Example 2.4 above for $f=Q_{1,0}^{[m]_{q}+1}$ and $g=0$. Thus when working with iterated delta operators, for example, $\delta_{2}^{2}(f)=\delta_{2}(\delta_{2}f)$, we cannot take $\delta_{2}f$ modulo $I_{m}$ before applying $\delta_{2}$ again.
@@ 2.4. The delta operator and the Dickson algebra
For each composition $\alpha$ of $n$, the invariant ring ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ is a module over the Dickson algebra ${\mathcal{D}}_{n}$. The next result describes how the delta operator interacts with the Dickson algebra in low ranks.
@@
We will make use frequently of the relation 2.1 relating upper-triangular invariants $V_{k}$ and Dickson invariants $Q_{n,i}$.
The first part is clear since $Q_{s,0}=L_{s}^{q-1}$. Note that the numerator of $\delta_{s}(f)$ always vanishes in ${\mathcal{Q}}_{m}$ by considering the Laplace expansion along the last row.
Proof of (2). Since $f\in\mathcal{D}_{1}$, let us assume that $f=Q_{1,0}^{s}$ for some $s\geq 0$. Since
$Q_{2,1}=V_{2}^{q-1}(x_{1},x_{2})+V_{1}^{q(q-1)}(x_{1})=V_{2}^{q-1}(x_{2},x_{1})+V_{1}^{q(q-1)}(x_{2}),$
the numerator of the difference $Q_{2,1}\delta_{2}(f)-\delta_{2}(Q_{1,0}^{q}f)$ has the form
$\begin{vmatrix}x_{1}&x_{2}\\ x_{1}^{q^{m}}V_{2}(x_{2},x_{1})^{q-1}x_{2}^{s(q-1)}&x_{2}^{q^{m}}V_{2}(x_{1},x_{2})^{q-1}x_{1}^{s(q-1)}\end{vmatrix},$
Observe that $x_{1}V_{2}(x_{1},x_{2})=x_{2}V_{2}(x_{2},x_{1})=L_{2}(x_{1},x_{2})$, it follows that
$Q_{2,1}\delta_{2}(f)-\delta_{2}(Q_{1,0}^{q}f)=x_{2}^{q^{m}}V_{2}(x_{1},x_{2})^{q-2}x_{1}^{s(q-1)}-x_{1}^{q^{m}}V_{2}(x_{2},x_{1})^{q-2}x_{2}^{s(q-1)}=0\quad\text{in ${\mathcal{Q}}_{m}(2)$}.$
@@
Proof of (3). We will treat the case $i=1$, as the situation for $i=2$ is similar. Consider the difference $Q_{3,1}\delta_{3}(f)-\delta_{3}(Q_{2,0}^{q}f)$. Using the identity $Q_{3,1}=V_{3}^{q-1}Q_{2,1}+Q_{2,0}^{q}$, the (3,1)-entry in the last row of the determinant in the numerator can be simplified as
$x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3})f(x_{2},x_{3}).$
There is a similar description for other entries in this row. Since
$V_{3}(x_{2},x_{3},x_{1})L_{2}(x_{2},x_{3})=L_{3}(x_{2},x_{3},x_{1})=L_{3}(x_{1},x_{2},x_{3}),$
@@
the Laplace expansion along the last row of the determinant in the numerator shows that the resulting polynomial belongs to $I_{m}$, which is zero in ${\mathcal{Q}}_{m}(3)$.
Proof of (4). From Proposition 2.1, we have $Q_{3,2}=V_{3}^{q-1}+Q_{2,1}^{q}$. We write $Q_{3,2}$ in two different ways:
$Q_{3,2}=V_{3}^{q-1}(x_{2},x_{3},x_{1})+Q_{2,1}^{q}(x_{2},x_{3})=V_{3}^{q-1}(x_{1},x_{3},x_{2})+Q_{2,1}^{q}(x_{1},x_{3}).$
It follows that $Q_{3,2}\delta_{2}(\delta_{2}f)-\delta_{2}(Q_{2,1}^{q}\delta_{2}f)$ equals
$\frac{1}{L_{2}(x_{1},x_{2})}\begin{vmatrix}x_{1}&x_{2}\\ x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})(\delta_{2}f)(x_{2},x_{3})&x_{2}^{q^{m}}V_{3}^{q-1}(x_{1},x_{3},x_{2})(\delta_{2}f)(x_{1},x_{3})\end{vmatrix}.$
Since $x_{2}V_{3}(x_{2},x_{3},x_{1})$ and $x_{1}V_{3}(x_{1},x_{3},x_{2})$ are divisible by $L_{2}(x_{1},x_{2})$, the above expression is a polynomial in the ideal generated by $(x_{1}^{q^{m}},x_{2}^{q^{m}})$. Thus
$Q_{3,2}\delta_{2}(\delta_{2}f)=\delta_{2}(Q_{2,1}^{q}\delta_{2}f)\quad\text{in${\mathcal{Q}}_{m}(3)$}.$
At this point, as noted in Warning 2.6, we cannot apply immediately part (2), replacing $Q_{2,1}^{q}\delta_{2}f$ with $\delta_{2}(Q_{1,0}^{q^{2}}f)$ on the right-hand side. Rather, we make use of the above strategy one more time, writing
$Q_{2,1}(x_{2},x_{3})=V_{2}^{q-1}(x_{3},x_{2})+V_{1}^{q^{2}(q-1)}(x_{3}),$
and similarly for $Q_{2,1}(x_{1},x_{3})$. The difference $\delta_{2}(Q_{2,1}^{q}\delta_{2}f)-\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ can be written as the sum
$x_{3}^{q^{m}}\frac{x_{2}^{q^{m}}x_{1}f(x_{1})V_{2}^{q^{2}-q-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}f(x_{2})V_{2}^{q^{2}-q-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}-\\ x_{2}^{q^{m}}f(x_{3})\frac{x_{1}x_{1}^{q^{m}}V_{2}^{q^{2}-q-1}(x_{3},x_{1})-x_{2}x_{1}^{q^{m}}V_{2}^{q^{2}-q-1}(x_{3},x_{2})}{L_{2}(x_{1},x_{2})}.$
Now the first summand is a polynomial (since the numerator is divisible by $x_{1}$ and vanishes whenever there is a linear relation $x_{2}=a_{1}x_{1}$) and is a multiple of $x_{3}^{q^{m}}$. Similarly, the second summand belongs to the ideal $(x_{2}^{q^{m}})$.
Proof of (5). We first apply the same strategy as part (4), writing
$Q_{3,1}=V_{3}^{q-1}(x_{1},x_{3},x_{2})Q_{2,1}(x_{1},x_{3})+Q_{2,0}^{q}(x_{3},x_{2})=V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3})+Q_{2,0}^{q}(x_{3},x_{1}),$
and obtain an equality in ${\mathcal{Q}}_{m}(3)$:
$Q_{3,1}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,0}^{q}\delta_{2}f).$
The right-hand side, by direct inspection, can be written as
$x_{3}^{q^{m}}\frac{x_{2}^{q^{m}}x_{1}^{2}f(x_{1})L_{2}^{q^{2}-q-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}^{2}f(x_{2})L_{2}^{q^{2}-q-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}+\\ x_{2}^{q^{m}}f(x_{3})x_{3}\frac{x_{1}x_{1}^{q^{m}}L_{2}^{q^{2}-q-1}(x_{1},x_{3})-x_{2}x_{1}^{q^{m}}L_{2}^{q^{2}-q-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}.$
By the same argument as in part (4), the first summand is divisible by $x_{3}^{q^{m}}$ and the second by $x_{2}^{q^{m}}$. ∎